# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@ 2.2. Other-Sensor-based Animal Pose Estimation Dataset
As discussed in section 2.1, researchers have traditionally relied on manual annotations or triangulation techniques to obtain 3D ground truth pose labels. An alternative approach for achieving high accuracy and low-noise 3D ground-truth animal pose data is to use a maker-based MoCap system with RGB cameras. For example, the Rat7M (2021) (Dunn et al., 2021) and PFERD (2024) (Li et al., 2024c) utilize MoCap system with RGB cameras and a set of makers to generate 3D ground truth pose data for rats and horses, respectively. Although these are effective for tracking individual animals from multiple views with temporal consistency, these datasets are limited by their focus on single animals and dependence on a large number of markers. To address these limitations, recent datasets like PAIR-R24M (2021) (Marshall et al., 2021) and 3D-POP (2023) (Naik et al., 2023) have been developed. These datasets use MoCap systems with fewer markers (12 and 8 markers, respectively) to track the 3D postures of multiple animals in the same frame, offering more scalable solutions for multi-animal tracking. Similarly, a recent dataset (Kearney et al., 2020a) utilizes MoCap systems with 63–82 markers to generate highly accurate 3D ground truth pose labels for dogs. This dataset also integrates depth sensors (i.e. RGBD), which enable us to develop more flexible solutions for 3D APE without relying on physical markers. Although the MoCap-based dataset provides high-quality 3D ground truth pose labels, APE algorithms trained using these MoCap datasets may exhibit poor generalization to in-the-wild scenarios, as the data collection for the MoCap-based datasets is all done indoors in a controlled environment, as you can see in table 4.
@@
WildPose (Muramatsu et al., 2024) is another promising multi-sensor system that combine RGB and LiDAR data. This system captures 2D RGB video and 3D point cloud data from a distance (up to 120 meters), making it particularly suited for studying free-ranging animals in natural habitats.
@@
In addition to these technologies, infrared cameras for camera-trapping in the wild, which use infrared LEDs to illuminate the scene to enable nighttime observations, have also been employed for wild animal pose labelling. For instance, the LoTE-Animal (2023) (Liu et al., 2023) uses infrared trap cameras to collect large-scale video footage of endangered animals across 11 species. This dataset captures a wide range of variations—such as ecological seasons, weather conditions, viewpoints, and habitats—and provides both instance segmentation labels and 2D pose keypoint annotations. The LoTE-Animal dataset is especially valuable for advancing semi-supervised and self-supervised APE models, with potential applications in conservation efforts for endangered species.
@@
Moreover, the VAREN (2024) dataset (Zuffi et al., 2024) focuses on 3D mesh representation, featuring 4k 3D scans of horses provided by the 3D scanner; this further advances 3D modelling and representation in APE research.
@@
Lastly, the IMU-based animal pose dataset collected by (Yigit et al., 2022) consists of acceleration and gyroscope measurements from sensors attached to the horse’s limbs. The data captures each limb’s 3D motion characteristics (like linear acceleration and angular velocity). It provides 3D information on movement and orientation but lacks absolute coordinates, as IMUs only capture relative motion.
@@
3. Error Metrics and Evaluation Methods
@@
3.1. RGB-based 2D Animal Pose Estimation Metric
@@
2D single-animal pose estimation methods like DeepLabCut (Mathis et al., 2018) and LEAP (Pereira et al., 2019) use a common metric like Root Means Square Error (RMSE) to evaluate their performance. RMSE is versatile and can be used for both 2D and 3D pose estimation and for evaluating various types of keypoints, regardless of the specific application domain. In addition to RMSE, many APE methods also use the Probability of Correct Keypoint (PCK), which measures the fraction of joints where the predicted joint positions are within a certain threshold distance from the ground truth joint positions. It is a normalised measure that indicates how well the keypoints are localised, and a higher PCK value indicates better localisation performance. Single-animal pose estimation methods like Synthetic-to-Real (Li and Lee, 2021b), Kim et al. (Kim et al., 2022), ScarceNet (Li and Lee, 2023), D-Gen (Li et al., 2023) and FSKD (Lu and Koniusz, 2022) use this metric for evaluation. Moreover, Li et al. (Li et al., 2020a) evaluate their single-animal APE methods with PCK, RMSE and Area Under Curve (AUC) along with an extension to these metrics with Permutation Invariant (PI). Li et al. (Li et al., 2020a) introduce PI metrics to evaluate the accuracy of pose estimation in challenging scenarios where it is difficult to uniquely identify limb identities (e.g. in fruit flies) with traditional metrics. PI metrics are designed to account for the ambiguity in identifying corresponding limbs or joints between the predicted pose and ground truth.
@@
Object Keypoint Similarity (OKS) has also been used in evaluating 2D single-animal pose estimation methods like D-Gen (Li et al., 2023); 2D multiple-animal pose estimation methods like DepthFormer (Liu et al., 2022b), Straka et al. (Straka et al., 2024), SLEAP (Pereira et al., 2022), and also 2D video based semi-supervised method like SemiMultiPose (Blau et al., 2022). OKS quantifies how well the keypoints are located to spatial scale and keypoint visibility, with closeness defined based on ground truth. The OKS lies in the range of 0 to 1, and a value close to 1 indicates that the detected keypoints coincide with the ground-truth keypoints. The OKS scores for every keypoint are computed for all objects in the dataset, and the average OKS is used to evaluate a model. Another common metric used for 2D APE is mean Average Precision (mAP), which has been used in evaluating ScarceNet (Li and Lee, 2023), SIPEC (Marks et al., 2022), SuperAnimal-Quadruped (Shaokai Ye and Mathis, 2024), OpenPose (Cao et al., 2017), DLCRNet (Lauer et al., 2022), maDLC (Lauer et al., 2022), SLEAP (Pereira et al., 2022), SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SemiMultiPose (Blau et al., 2022). The mAP is an evaluation metric based on OKS and has become widely used for pose estimation, particularly following its adoption in the COCO keypoint challenge. OKS measures the similarity between the predicted keypoints and the ground truth, and Average Precision (AP) is then calculated to assess the prediction’s performance on a dataset by comparing it against different OKS thresholds. The mAP is the mean of several AP values, each calculated using a different threshold.
@@
Last but not least, another metric that has been used for the 2D video-based APE method evaluation is Canonical Correlation Analysis (CCA), which is a statistical technique used to examine the relationship between two sets of variables. It determines linear combinations of the variables in each set that are maximally correlated; this metric has been used for evaluating the Lightning Pose (Biderman et al., 2023) method, however, we have not seen wide adoption of this evaluation metric within the APE community.
3.2. RGB-based 3D Animal Pose Estimation Metric
@@
3D APE metrics can also incorporate some of the metrics used in 2D APE. For example, PCK is employed in methods such as Three-D Safari (Zuffi et al., 2019a), SMBLD (Biggs et al., 2020), Li et al. (Li and Lee, 2021a), Birds of a Feather (Wang et al., 2021a), BITE (Rüegg et al., 2023), MagicPony (Wu et al., 2023b), DigiDogs (Shooter et al., 2024), Monet (Yao et al., 2019), DAN-NCE (Dunn et al., 2021), and Muramatsu et al. (Muramatsu et al., 2022). The well-known RMSE metric is also used in 3D APE methods such as Hu et al. (Hu et al., 2023), Muramatsu et al. (Joska et al., 2021), 3D-UPPER (Ebrahimi et al., 2023), and SBeA (Han et al., 2024).
@@
Apart from PCK and RMSE, 3D pose estimation-specific metrics such as Mean Per Joint Position Error (MPJPE) have been used in 3D APE methods like DigiDogs (Shooter et al., 2024) and MAMMAL (An et al., 2023). MPJPE measures the average Euclidean distance between the predicted and ground truth joint positions across all joints in the skeleton, quantifying the accuracy of the joint position estimation. A lower MPJPE indicates that the predicted joint positions are closer to the actual joint locations. An extension of the MPJPE metric is N-MPJPE, which has been used by Dai et al. (Dai et al., 2023) for their 3D APE model. N-MPJPE involves normalising the ground truth 3D pose by subtracting the mean and dividing by the standard deviation, aligning the predicted 3D pose with the normalised ground truth using Procrustes alignment, and then calculating the MPJPE, which results in Normalized-MPJPE.
@@
For mesh-based APE as discussed in section 4.3.1, a well-known evaluation metric is the 3D Chamfer Distance, which has been used in several model-free methods such as DOVE (Wu et al., 2023a), MagicPony (Wu et al., 2023b), and BANMo (Yang et al., 2022a). The 3D Chamfer Distance measures the average distance between the reconstructed surface points and the ground truth surface points. It finds the nearest neighbour matches between the two sets of points and computes the distance, which is sensitive to outliers. A lower value indicates better reconstruction quality.
@@
Another approach in mesh-based APE is the model-based pose estimation methods, for example, SMAL (Zuffi et al., 2017). In these approaches, metrics like Scan-to-Mesh distance, keypoint reprojection error, and silhouette reprojection error are commonly used. The Scan-to-Mesh distance measures the accuracy of the reconstructed 3D shape by calculating the distance from each point on the scan to the closest point on the predicted mesh surface. Keypoint reprojection error evaluates the pose estimation by comparing the 2D projection of 3D keypoints from the predicted model with the annotated 2D keypoints in the images. It assesses how well the 3D pose model aligns with the 2D image data. Similarly, the silhouette reprojection error measures how accurately the 3D model silhouette matches the ground truth silhouette in the 2D image, indicating how well the 3D shape projection aligns with the visible boundaries of the animal in the image. Other model-based methods, such as SMBLD (Biggs et al., 2020) and Li et al. (Li and Lee, 2021a), use metrics like PCK and Intersection over Union (IoU) for performance evaluation.
@@
Moreover, due to the scarcity of animal datasets that include 3D pose annotations, comprehensive quantitative evaluations of 3D pose predictions are not always feasible. Therefore, researchers often assess 3D pose predictions using qualitative means, as Sosa et al. (Sosa and Hogg, 2023) explained. Examples of qualitative evaluations can be found in Salem et al. (Salem et al., 2019), Sosa et al. (Sosa and Hogg, 2023), Monet (Yao et al., 2019), and Muramatsu et al. (Joska et al., 2021).
3.3. Other-Modality-based Animal Pose Estimation Metric
@@
Different sensor modalities have different ways of evaluating their APE models. For example, an IMU-based APE method (Yigit et al., 2022) is evaluated by calculating the mean and standard deviation (Std) of the estimated joint angle errors and comparing the predictions with ground truth data obtained through a MoCap system. Additionally, the model is assessed for Intra-individual Variability (IAV) and Inter-individual Variability (IEV), which quantify the consistency of limb movement across different gait cycles and between different horses. Another metric used is the Variance Ratio (VR), which captures the repeatability of the gait cycle by evaluating how similar joint movements are within multiple cycles. This thorough approach ensures that this IMU-based model’s performance is evaluated for accuracy and consistency across different conditions.
@@
On the other hand, one of the RGBD-based APE methods, RGBD-Dog (Kearney et al., 2020a), is evaluated using MPJPE and PCK. The evaluation is conducted across different body parts and for entire dog poses. Additionally, Procrustes Analysis (PA) is applied to align the predicted skeleton with the ground truth skeleton. The PA is a statistical shape analysis technique. In the context of pose estimation, it aligns the predicted 3D skeleton with the ground truth skeleton by minimising the sum of squared distances between corresponding points in the two sets. In the RGBD-Dog (Kearney et al., 2020a), PA is applied to refine the existing PCK and MPJPE evaluation metrics to PA-PCK, PA-PCK 3D and PA-MPJPE, which is then used to measure the error after aligning the predicted and ground truth skeletons to focus purely on pose accuracy. Other multi-modal systems, for example, Sound of Motion (RGB+acoustic) (Li et al., 2022) and CLHOP (audio+video) (Li et al., 2024a), also use PA-MPJPE for their APE evaluation in addition to PCK and IoU, which are common in the RGB-based APE evaluations.
@@
4. RGB-based Animal Pose Estimation
@@
In recent years, APE has achieved significant progress. Among the various modalities utilised for APE, RGB-based approaches have emerged as the most developed and extensively studied research method (Pereira et al., 2019; Cao et al., 2019; Li and Lee, 2021b; Lauer et al., 2022). These approaches leverage visual information captured by RGB cameras, take animal images or videos as inputs, and subsequently produce the animal’s pose, which manifests as a 2D skeletal keypoints representation of pose, 3D skeletal keypoints representation of pose, or 3D animal mesh reconstructions, as exemplified in LABEL:fig:Skeleton_Pose_Representations and LABEL:fig:Body_Mesh_Reconstructions. In addition to skeletal keypoints and mesh representation for pose estimations, dense pose representations have also been used mainly for humans (Riza Alp Güler, 2018) and chimpanzees (Sanakoyeu et al., 2020). In this paper, skeletal keypoints and mesh representations APE have been mainly analysed in detail, while keypoint-based methods remain dominant due to their established presence, easier access to annotated data, and lower computational requirements.
@@
Before we go into the details of RGB-based APE methods, let us first discuss the development of this branch. Traditionally, we use direct visual observation to extract the animal poses, which can offer real-time and context-rich data. However, it suffers from inherent drawbacks, including observer bias, mental fatigue (Graving et al., 2019), and long data collection and observation times. Later, researchers start to use video monitoring and recording to expedite data collection and minimise the influence of human involvement (Graving et al., 2019). However, manual scoring of videos by eyes still remains time-consuming and often fails to capture subtle and fast-paced movements that are integral to various behaviours. Moreover, the reliance on human observers restricts the dataset size, this may also lead to statistical errors that limit researchers’ ability to address scientific questions accurately.
@@
Recently, automated body kinematics tracking techniques utilising advanced imaging software and hardware have been developed to tackle the challenges in manual scoring. Marker-based methods rely on attaching markers (e.g. MoCap suits (Kearney et al., 2020b), animal-borne cameras (Patel et al., 2017)) to animals for posture measurements. While these methods can achieve precise pose estimation, the reliance on physical markers can pose practical challenges, especially when investigating innate behaviours in natural habitats (Wu et al., 2023b) or when observing the movements of small creatures like mice (Lauer et al., 2022) and fish. Furthermore, most marker-based approaches are limited to highly controlled lab environments and specialised hardware setups. Such methodologies are often impractical when the target animals demonstrate varied behaviours within unfamiliar environments.
@@
With the rapid advancement in Computer Vision (CV), many automated markerless APE methods (Karashchuk et al., 2021; Mathis et al., 2018; Lauer et al., 2022; Pereira et al., 2019; Graving et al., 2019) start to employ supervised learning to directly infer and estimate animals’ pose and movements from images or videos, removing the need for physical markers that may alter natural behaviour. For example, the markerless method DeepLabCut (Mathis et al., 2018) has been used for behaviour tracking in wild chimpanzees and bonobos (Wiltshire et al., 2023).
@@
However, a significant obstacle in APE is the acquisition of accurate ground truth poses as discussed in section 2. To address the scarcity of large-scale, well-labelled animal pose datasets, many weakly and semi-supervised (Cao et al., 2019), semi-supervised (Kim et al., 2022; Blau et al., 2022; Biderman et al., 2023; Yao et al., 2019), self-supervised (Sosa and Hogg, 2023; Shooter et al., 2024; Yang et al., 2022a; Han et al., 2024) and unsupervised learning (Dai et al., 2023; Kim et al., 2022; Li and Lee, 2021b; Ebrahimi et al., 2023; Shaokai Ye and Mathis, 2024; Wu et al., 2023a, b; Li et al., 2024b) approaches have been developed. For instance, recent literature (Dai et al., 2023) proposes an unsupervised 3D animal canonical pose estimation method with geometric self-supervision; this method necessitates only 2D poses for training, which significantly alleviates the requirement for extensively labelled data. Current APE research trends show growing interest in unsupervised, semi-supervised, and self-supervised methods due to the mentioned challenges to reach the same performance level as supervised methods. However, supervised methods are still a significant focus in current APE studies, especially with the emergence of larger datasets and benchmarking platforms like AP-10K (Yu et al., 2021b), APT-36K (Yang et al., 2022b), TigDog (Del Pero et al., 2015, 2017), and AcinoSet (Joska et al., 2021), as we can see in Table 5-Table 7.
@@
4.1. 2D Animal Pose Estimation from an Image
@@
This section focuses on 2D APE methods that do not use temporal information to localize the 2D coordinates of an animal’s or animal’s joints from a single image frame (table 5). Note that the majority of the APE algorithms that can extract poses from videos (i.e. multiple frames) can also extract poses from a single image. In Table 5, column ”I,V” provides the information on APE methods that can be applied on both an image and videos; or only on an image.
@@
Based on the number of target animals being processed in the scene, this pose estimation problem can be further divided into single-animal pose estimation (section 4.1.1) and multiple-animal pose estimation (section 4.1.2). While pose estimation in individual animals has received significant attention and serves as the foundation for most APE studies, extending it to multiple-animal pose estimation is an active area of research that is still evolving (Pereira et al., 2022).
4.1.1. Single Animal

### v2

@@ 2.2 Other-Sensor-based Animal Pose Estimation Dataset
As discussed in Section 2.1, researchers have traditionally relied on manual annotations or triangulation techniques to obtain 3D ground truth pose labels. An alternative approach for achieving high accuracy and low-noise 3D ground-truth animal pose data is to use a maker-based MoCap system with RGB cameras. For example, the Rat7M (2021) [30] and PFERD (2024) [65] utilize MoCap system with RGB cameras and a set of makers to generate 3D ground truth pose data for rats and horses, respectively. Although these are effective for tracking individual animals from multiple views with temporal consistency, these datasets are limited by their focus on single animals and dependence on a large number of markers. To address these limitations, recent datasets like PAIR-R24M (2021) [89] and 3D-POP (2023) [99] have been developed. These datasets use MoCap systems with fewer markers (12 and 8 markers, respectively) to track the 3D postures of multiple animals in the same frame, offering more scalable solutions for multi-animal tracking. Similarly, a recent dataset [56] utilizes MoCap systems with 63–82 markers to generate highly accurate 3D ground truth pose labels for dogs. This dataset also integrates depth sensors (i.e. RGBD), which enable us to develop more flexible solutions for 3D APE without relying on physical markers. Although the MoCap-based dataset provides high-quality 3D ground truth pose labels, APE algorithms trained using these MoCap datasets may exhibit poor generalization to in-the-wild scenarios, as the data collection for the MoCap-based datasets is all done indoors in a controlled environment, as you can see in Table 4.
@@
WildPose [97] is another promising multi-sensor system that combine RGB and LiDAR data. This system captures 2D RGB video and 3D point cloud data from a distance (up to 120 meters), making it particularly suited for studying free-ranging animals in natural habitats.
@@
In addition to these technologies, infrared cameras for camera-trapping in the wild, which use infrared LEDs to illuminate the scene to enable nighttime observations, have also been employed for wild animal pose labelling. For instance, the LoTE-Animal (2023) [77] uses infrared trap cameras to collect large-scale video footage of endangered animals across 11 species. This dataset captures a wide range of variations—such as ecological seasons, weather conditions, viewpoints, and habitats—and provides both instance segmentation labels and 2D pose keypoint annotations. The LoTE-Animal dataset is especially valuable for advancing semi-supervised and self-supervised APE models, with potential applications in conservation efforts for endangered species.
@@
Moreover, the VAREN (2024) dataset [174] focuses on 3D mesh representation, featuring 4k 3D scans of horses provided by the 3D scanner; this further advances 3D modelling and representation in APE research.
@@
Lastly, the IMU-based animal pose dataset collected by [162] consists of acceleration and gyroscope measurements from sensors attached to the horse’s limbs. The data captures each limb’s 3D motion characteristics (like linear acceleration and angular velocity). It provides 3D information on movement and orientation but lacks absolute coordinates, as IMUs only capture relative motion.
@@
3 Error Metrics and Evaluation Methods
@@
3.1 RGB-based 2D Animal Pose Estimation Metric
@@
2D single-animal pose estimation methods like DeepLabCut [90] and LEAP [111] use a common metric like Root Means Square Error (RMSE) to evaluate their performance. RMSE is versatile and can be used for both 2D and 3D pose estimation and for evaluating various types of keypoints, regardless of the specific application domain. In addition to RMSE, many APE methods also use the Probability of Correct Keypoint (PCK), which measures the fraction of joints where the predicted joint positions are within a certain threshold distance from the ground truth joint positions. It is a normalised measure that indicates how well the keypoints are localised, and a higher PCK value indicates better localisation performance. Single-animal pose estimation methods like Synthetic-to-Real [69], Kim et al. [60], ScarceNet [70], D-Gen [71] and FSKD [83] use this metric for evaluation. Moreover, Li et al. [72] evaluate their single-animal APE methods with PCK, RMSE and Area Under Curve (AUC) along with an extension to these metrics with Permutation Invariant (PI). Li et al. [72] introduce PI metrics to evaluate the accuracy of pose estimation in challenging scenarios where it is difficult to uniquely identify limb identities (e.g. in fruit flies) with traditional metrics. PI metrics are designed to account for the ambiguity in identifying corresponding limbs or joints between the predicted pose and ground truth.
@@
Object Keypoint Similarity (OKS) has also been used in evaluating 2D single-animal pose estimation methods like D-Gen [71]; 2D multiple-animal pose estimation methods like DepthFormer [78], Straka et al. [134], SLEAP [112], and also 2D video based semi-supervised method like SemiMultiPose [12]. OKS quantifies how well the keypoints are located to spatial scale and keypoint visibility, with closeness defined based on ground truth. The OKS lies in the range of 0 to 1, and a value close to 1 indicates that the detected keypoints coincide with the ground-truth keypoints. The OKS scores for every keypoint are computed for all objects in the dataset, and the average OKS is used to evaluate a model. Another common metric used for 2D APE is mean Average Precision (mAP), which has been used in evaluating ScarceNet [70], SIPEC [88], SuperAnimal-Quadruped [160], OpenPose [16], DLCRNet [64], maDLC [64], SLEAP [112], SuperAnimal-TopViewMouse [160] and SemiMultiPose [12]. The mAP is an evaluation metric based on OKS and has become widely used for pose estimation, particularly following its adoption in the COCO keypoint challenge. OKS measures the similarity between the predicted keypoints and the ground truth, and Average Precision (AP) is then calculated to assess the prediction’s performance on a dataset by comparing it against different OKS thresholds. The mAP is the mean of several AP values, each calculated using a different threshold.
@@
Last but not least, another metric that has been used for the 2D video-based APE method evaluation is Canonical Correlation Analysis (CCA), which is a statistical technique used to examine the relationship between two sets of variables. It determines linear combinations of the variables in each set that are maximally correlated; this metric has been used for evaluating the Lightning Pose [9] method, however, we have not seen wide adoption of this evaluation metric within the APE community.
3.2 RGB-based 3D Animal Pose Estimation Metric
@@
3D APE metrics can also incorporate some of the metrics used in 2D APE. For example, PCK is employed in methods such as Three-D Safari [176], SMBLD [10], Li et al. [68], Birds of a Feather [144], BITE [120], MagicPony [148], DigiDogs [128], Monet [159], DAN-NCE [30], and Muramatsu et al. [96]. The well-known RMSE metric is also used in 3D APE methods such as Hu et al. [47], Muramatsu et al. [53], 3D-UPPER [31], and SBeA [43].
@@
Apart from PCK and RMSE, 3D pose estimation-specific metrics such as Mean Per Joint Position Error (MPJPE) have been used in 3D APE methods like DigiDogs [128] and MAMMAL [2]. MPJPE measures the average Euclidean distance between the predicted and ground truth joint positions across all joints in the skeleton, quantifying the accuracy of the joint position estimation. A lower MPJPE indicates that the predicted joint positions are closer to the actual joint locations. An extension of the MPJPE metric is N-MPJPE, which has been used by Dai et al. [19] for their 3D APE model. N-MPJPE involves normalising the ground truth 3D pose by subtracting the mean and dividing by the standard deviation, aligning the predicted 3D pose with the normalised ground truth using Procrustes alignment, and then calculating the MPJPE, which results in Normalized-MPJPE.
@@
For mesh-based APE, a well-known evaluation metric is the 3D Chamfer Distance, which has been used in several model-free methods such as DOVE [147], MagicPony [148], and BANMo [153]. The 3D Chamfer Distance measures the average distance between the reconstructed surface points and the ground truth surface points. It finds the nearest neighbour matches between the two sets of points and computes the distance, which is sensitive to outliers. A lower value indicates better reconstruction quality.
@@
Another approach in mesh-based APE is the model-based pose estimation methods, for example, SMAL [178]. In these approaches, metrics like Scan-to-Mesh distance, keypoint reprojection error, and silhouette reprojection error are commonly used. The Scan-to-Mesh distance measures the accuracy of the reconstructed 3D shape by calculating the distance from each point on the scan to the closest point on the predicted mesh surface. Keypoint reprojection error evaluates the pose estimation by comparing the 2D projection of 3D keypoints from the predicted model with the annotated 2D keypoints in the images. It assesses how well the 3D pose model aligns with the 2D image data. Similarly, the silhouette reprojection error measures how accurately the 3D model silhouette matches the ground truth silhouette in the 2D image, indicating how well the 3D shape projection aligns with the visible boundaries of the animal in the image. Other model-based methods, such as SMBLD [10] and Li et al. [68], use metrics like PCK and Intersection over Union (IoU) for performance evaluation.
@@
Moreover, due to the scarcity of animal datasets that include 3D pose annotations, comprehensive quantitative evaluations of 3D pose predictions are not always feasible. Therefore, researchers often assess 3D pose predictions using qualitative means, as Sosa et al. [132] explained. Examples of qualitative evaluations can be found in Salem et al. [123], Sosa et al. [132], Monet [159], and Muramatsu et al. [53].
3.3 Other-Modality-based Animal Pose Estimation Metric
@@
Different sensor modalities have different ways of evaluating their APE models. For example, an IMU-based APE method [162] is evaluated by calculating the mean and standard deviation (Std) of the estimated joint angle errors and comparing the predictions with ground truth data obtained through a MoCap system. Additionally, the model is assessed for Intra-individual Variability (IAV) and Inter-individual Variability (IEV), which quantify the consistency of limb movement across different gait cycles and between different horses. Another metric used is the Variance Ratio (VR), which captures the repeatability of the gait cycle by evaluating how similar joint movements are within multiple cycles. This thorough approach ensures that this IMU-based model’s performance is evaluated for accuracy and consistency across different conditions.
@@
On the other hand, one of the RGBD-based APE methods, RGBD-Dog [56], is evaluated using MPJPE and PCK. The evaluation is conducted across different body parts and for entire dog poses. Additionally, Procrustes Analysis (PA) is applied to align the predicted skeleton with the ground truth skeleton. The PA is a statistical shape analysis technique. In the context of pose estimation, it aligns the predicted 3D skeleton with the ground truth skeleton by minimising the sum of squared distances between corresponding points in the two sets. In the RGBD-Dog [56], PA is applied to refine the existing PCK and MPJPE evaluation metrics to PA-PCK, PA-PCK 3D and PA-MPJPE, which is then used to measure the error after aligning the predicted and ground truth skeletons to focus purely on pose accuracy. Other multi-modal systems, for example, Sound of Motion (RGB+acoustic) [67] and CLHOP (audio+video) [66], also use PA-MPJPE for their APE evaluation in addition to PCK and IoU, which are common in the RGB-based APE evaluations.
@@
4 RGB-based Animal Pose Estimation
@@
In recent years, APE has achieved significant progress. Among the various modalities utilised for APE, RGB-based approaches have emerged as the most developed and extensively studied research method [111, 15, 69, 64]. These approaches leverage visual information captured by RGB cameras, take animal images or videos as inputs, and subsequently produce the animal’s pose, which manifests as a 2D skeletal keypoints representation of pose, 3D skeletal keypoints representation of pose, or 3D animal mesh reconstructions, as exemplified in Figure 1(a) and Figure 1(b). In addition to skeletal keypoints and mesh representation for pose estimations, dense pose representations have also been used mainly for humans [118] and chimpanzees [124]. In this paper, skeletal keypoints and mesh representations APE have been mainly analysed in detail, while keypoint-based methods remain dominant due to their established presence, easier access to annotated data, and lower computational requirements.
@@
Before we go into the details of RGB-based APE methods, let us first discuss the development of this branch. Traditionally, we use direct visual observation to extract the animal poses, which can offer real-time and context-rich data. However, it suffers from inherent drawbacks, including observer bias, mental fatigue [40], and long data collection and observation times. Later, researchers start to use video monitoring and recording to expedite data collection and minimise the influence of human involvement [40]. However, manual scoring of videos by eyes still remains time-consuming and often fails to capture subtle and fast-paced movements that are integral to various behaviours. Moreover, the reliance on human observers restricts the dataset size, this may also lead to statistical errors that limit researchers’ ability to address scientific questions accurately.
@@
Recently, automated body kinematics tracking techniques utilising advanced imaging software and hardware have been developed to tackle the challenges in manual scoring. Marker-based methods rely on attaching markers (e.g. MoCap suits [56], animal-borne cameras [108]) to animals for posture measurements. While these methods can achieve precise pose estimation, the reliance on physical markers can pose practical challenges, especially when investigating innate behaviours in natural habitats [148] or when observing the movements of small creatures like mice [64] and fish. Furthermore, most marker-based approaches are limited to highly controlled lab environments and specialised hardware setups. Such methodologies are often impractical when the target animals demonstrate varied behaviours within unfamiliar environments.
@@
With the rapid advancement in Computer Vision (CV), many automated markerless APE methods [55, 90, 64, 111, 40] start to employ supervised learning to directly infer and estimate animals’ pose and movements from images or videos, removing the need for physical markers that may alter natural behaviour. For example, the markerless method DeepLabCut [90] has been used for behaviour tracking in wild chimpanzees and bonobos [146].
@@
However, a significant obstacle in APE is the acquisition of accurate ground truth poses as discussed in Section 2. To address the scarcity of large-scale, well-labelled animal pose datasets, many weakly and semi-supervised [15], semi-supervised [60, 12, 9, 159], self-supervised [132, 128, 153, 43] and unsupervised learning [19, 60, 69, 147, 31, 148, 160, 74] approaches have been developed. For instance, recent literature [19] proposes an unsupervised 3D animal canonical pose estimation method with geometric self-supervision; this method necessitates only 2D poses for training, which significantly alleviates the requirement for extensively labelled data. Current APE research trends show growing interest in unsupervised, semi-supervised, and self-supervised methods due to the mentioned challenges to reach the same performance level as supervised methods. However, supervised methods are still a significant focus in current APE studies, especially with the emergence of larger datasets and benchmarking platforms like AP-10K [164], APT-36K [157], TigDog [22, 23], and AcinoSet [53], as we can see in Table 5-Table 7.
@@
4.1 2D Animal Pose Estimation from an Image
@@
This section focuses on 2D APE methods that do not use temporal information to localize the 2D coordinates of an animal’s or animal’s joints from a single image frame (Table 5). Note that the majority of the APE algorithms that can extract poses from videos (i.e. multiple frames) can also extract poses from a single image. In Table 5, column “I,V” provides the information on APE methods that can be applied on both an image and videos; or only on an image.
@@
Based on the number of target animals being processed in the scene, this pose estimation problem can be further divided into single-animal pose estimation (Section 4.1.1) and multiple-animal pose estimation (Section 4.1.2). While pose estimation in individual animals has received significant attention and serves as the foundation for most APE studies, extending it to multiple-animal pose estimation is an active area of research that is still evolving [112].
4.1.1 Single Animal